# E-MoFlow: Learning Egomotion and Optical Flow from Event Data via Implicit Regularization

### v1

@@ Abstract
The estimation of optical flow and 6-DoF ego-motion—two fundamental tasks in 3D vision—has typically been addressed independently. For neuromorphic vision (e.g., event cameras), however, the lack of robust data association makes solving the two problems separately an ill-posed challenge, especially in the absence of supervision via ground truth. Existing works mitigate this ill-posedness by either enforcing the smoothness of the flow field via an explicit variational regularizer or leveraging explicit structure-and-motion priors in the parametrization to improve event alignment. The former notably introduces bias in results and computational overhead, while the latter—which parametrizes the optical flow in terms of the scene depth and the camera motion—often converges to suboptimal local minima. To address these issues, we propose an unsupervised framework that jointly optimizes egomotion and optical flow via implicit spatial-temporal and geometric regularization. First, by modeling camera’s egomotion as a continuous spline and optical flow as an implicit neural representation, our method inherently embeds spatial-temporal coherence through inductive biases. Second, we incorporate structure-and-motion priors through differential geometric constraints, bypassing explicit depth estimation while maintaining rigorous geometric consistency. As a result, our framework (called E-MoFlow) unifies egomotion and optical flow estimation via implicit regularization under a fully unsupervised paradigm. Experiments demonstrate its versatility to general 6-DoF motion scenarios, achieving state-of-the-art performance among unsupervised methods and competitive even with supervised approaches.
@@ 2 Related Work
2.1 Event-based Optical Flow Learning
Event cameras measure per-pixel intensity changes asynchronously rather than capturing absolute intensity values at fixed intervals [12]. While this property enables high dynamic range and low-latency sensing, estimating optical flow from purely local event patches inherently suffers from the aperture problem [13, 14]. Since traditional optimization-based methods struggle to recover full motion fields from ambiguous local event data, learning-based approaches have emerged as the dominant solution. Existing learning-based optical flow methods can be broadly categorized into supervised, semi-supervised, and unsupervised approaches.
@@ 2.1 Event-based Optical Flow Learning
2.2 Event Camera Egomotion Estimation
@@
2.3 Flow and Motion Joint Estimation
@@ 4.1 Experimental Setup
We conduct comprehensive evaluations on the MVSEC dataset [53], which is the de facto standard dataset used in prior work to benchmark optical flow and 6-DoF ego-motion estimations. This dataset contains both indoor sequences recorded by drones and outdoor sequences recorded by vehicles, providing synchronized event streams, grayscale frames, along with ground truth optical flow and 6-DoF motion parameters. The event camera has a resolution of 346×260 pixels. In total, we evaluated on 63.5 million events spanning 265 seconds. For optical flow evaluation, we compute two standard metrics: endpoint error (EPE), and the percentage of pixels with EPE >> 3 pixels (denoted by "% Out"), exclusively over valid ground truth regions with event activity in the evaluation interval. For motion estimation accuracy, we adopt the root mean square error (RMSE) metric proposed in [14] to measure angular velocity ($^{\circ}/s$) and linear velocity ($m/s$) errors.
@@
Optical flow estimation using event camera is a widely explored task with various methodological paradigms. We categorize existing approaches into four primary classes: Supervised Learning(SL), which requires ground truch optical flow for training (e.g., E-RAFT [26], EV-FlowNet-EST [25], EV-FlowNet+ [29], DCEIFlow [54], TMA [24], ADM-Flow [31]); Semi-Supervised Learning (SSL), leveraging grayscale images as supervision through photometric loss construction (e.g., EV-FlowNet [18], STE-FlowNet [33]); Unsupervised Learning (USL), relying solely on event data by warping events to reference time and maximizing accumulated image contrast (e.g., MotionPriorCMax [36], ConvGRU-EV-FlowNet [32], FireFlowNet [27], USL-EV-FlowNet [19], ET-FlowNeT [35], EV-MGRFlowNet [55], Paredes et al. [37]); and Model-Based (MB) methods, which also adopt contrast maximization objectives but employ traditional nonlinear optimization instead of neural networks (e.g., MultiCM [17], Akolkar et al. [13], Nagata et al. [56], Brebion et al. [57], Cuadrado et al. [28], Shiba et al. [58]). From these paradigms, we select representative prior works as strong baselines for comprehensive quantitative benchmarking of optical flow estimation performance.
For 6-DoF motion estimation, existing methods can be categorized into two classes based on whether depth prior knowledge is required. Approaches such as AEmin [42], Incmin [59], and PEME [43] require depth-augmented event data to achieve a 6-DoF estimation, while ECN [48] and MultiCM [17] can perform a 6-DoF estimation relying solely on raw event streams without depth priors.
@@
We implemented our method with PyTorch. The neural implicit optical flow field network adopts the MLP architecture. The detailed network architecture can be found in our supplementary material. Following prior works [17] [43], we partition the entire sequence into multiple segments during training, each containing 30k events. Because the time interval of each segment is short, the cubic spline modeling continuous camera motion employs only 4 control knots, whose 6-dimensional vectors are initialized to a constant value of 0.2. To solve the event warping trajectory Eq. (2), we employ the euler solver for its computational efficiency in numerical integration. The weight of the differential geometric loss is set to 0.25 and the differential flow loss to 1. We use two separate Adam optimizers [60] for the neural implicit flow field $\text{NN}_{\theta}$ and camera motion $\beta$. The learning rate for the flow field is exponentially decayed from $1\times 10^{-4}$ to $6.3\times 10^{-5}$, while that for camera motion remains fixed at $1\times 10^{-3}$. Each short event segment is trained for 1k iterations. All experiments were conducted on a NVIDIA RTX 4090.
Table 1: Ablation studies on differential geometric constrain. We conducted ablation studies under the $dt=4$ settings with three configurations: a) w/o geometric constraints, b) w/ geometric constraints, and c) direct supervision using ground truth motion.
Table 2: Quantitative comparison of optical flow estimation task on MVSEC dataset. $dt=1$ denotes evaluating optical flow errors by a one-frame interval duration , while $dt=4$ indicates a four-frame interval duration. Bold is the best among all methods; underlined is second best.
indoor_flying1 indoor_flying2 indoor_flying3 outdoor_day1 average $dt=1$ EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ SL EV-FlowNet-EST [25] 0.97 0.91 1.38 8.20 1.43 6.47 EV-FlowNet+ [29] 0.56 1.00 0.66 1.00 0.59 1.00 0.68 0.99 0.623 0.998 E-RAFT [26] 1.10 5.72 1.94 30.79 1.66 25.20 0.24 1.70 1.235 15.853 DCEIFlow [54] 0.75 1.55 0.90 2.10 0.80 1.77 0.22 0.00 0.668 1.355 TMA [24] 1.06 3.63 1.81 27.29 1.58 23.26 0.25 0.07 1.175 13.563 ADM-Flow [31] 0.52 0.14 0.68 1.18 0.52 0.04 0.41 0.00 0.533 0.340 SSL EV-FlowNet [18] 1.03 2.20 1.72 15.10 1.53 11.90 0.49 0.20 1.193 7.350 STE-FlowNet [33] 0.57 0.10 0.79 1.60 0.72 1.30 0.42 0.00 0.625 0.750 MB Akolkar et al. [13] 1.52 1.59 1.89 2.75 1.938 Nagata et al. [56] 0.62 0.93 0.84 0.77 0.790 Brebion et al. [57] 0.52 0.10 0.98 5.50 0.71 2.10 0.53 0.20 0.685 1.975 Cuadrado et al. [28] 0.58 0.72 0.67 0.85 0.705 Shiba et al. [58] 1.05 2.90 1.68 13.44 1.43 8.97 0.94 3.08 1.275 7.098 MultiCM [17] 0.42 0.09 0.60 0.59 0.50 0.29 0.30 0.11 0.455 0.270 USL USL-EV-FlowNet [19] 0.58 0.00 1.02 4.00 0.87 3.00 0.32 0.00 0.698 1.750 FireFlowNet [27] 0.97 2.60 1.67 15.30 1.43 11.00 1.06 6.60 1.283 8.875 ConvGRU-EV-FlowNet [32] 0.60 0.51 1.17 8.06 0.93 5.64 0.47 0.25 0.793 3.615 ET-FlowNeT [35] 0.57 0.53 1.20 8.48 0.95 5.73 0.39 0.12 0.778 3.715 EV-MGRFlowNet [55] 0.41 0.17 0.70 2.35 0.59 1.29 0.28 0.02 0.495 0.958 Paredes et al. [37] 0.44 0.00 0.88 4.51 0.70 2.41 0.27 0.05 0.573 1.743 MotionPriorCMax [36] 0.45 0.09 0.71 2.40 0.60 0.93 Ours 0.40 0.30 0.52 0.18 0.46 0.29 0.42 0.54 0.450 0.328 $dt=4$ SL E-RAFT [26] 2.81 40.25 5.09 64.19 4.46 57.11 0.72 1.12 3.270 40.668 DCEIFlow [54] 2.08 21.47 3.48 42.05 2.51 29.73 0.89 3.19 2.240 24.110 TMA [24] 2.43 29.91 4.32 52.74 3.60 42.02 0.70 1.08 2.762 31.438 ADM-Flow [31] 1.42 7.78 1.88 16.70 1.61 11.40 1.51 10.20 1.605 11.520 SSL EV-FlowNet [18] 2.25 24.70 4.05 45.30 3.45 39.70 1.23 7.30 2.745 29.250 STE-FlowNet [33] 1.77 14.70 2.52 26.10 2.23 22.10 0.99 3.90 1.878 16.700 MB Shiba et al. [58] 4.06 53.88 6.39 71.82 5.36 65.57 3.60 49.04 4.853 60.077 MultiCM [17] 1.68 12.79 2.49 26.31 2.06 18.93 1.25 9.19 1.870 16.805 USL USL-EV-FlowNet [19] 2.18 24.20 3.85 46.80 3.18 47.80 1.30 9.70 2.628 32.125 ConvGRU-EV-FlowNet [32] 2.16 21.51 3.90 40.72 3.00 29.60 1.69 12.50 2.688 26.083 ET-FlowNeT [35] 2.08 20.02 3.99 41.33 3.13 31.70 1.47 9.17 2.667 25.555 EV-MGRFlowNet [55] 1.50 8.67 2.39 23.70 2.06 18.00 1.10 6.22 1.763 14.147 Ours 1.56 9.13 2.04 18.43 1.84 13.57 1.61 14.25 1.763 13.845
4.2 Ablation Study
To validate the efficacy of differential geometric constraints Eq. (9), we conducted ablation experiments on the MVSEC dataset [53] with three configurations: a) no geometric constraints, b) with differential geometric constraints, and c) direct supervised by ground truth motion. As demonstrated in Tab. 1, the differential geometric constraints yield improvements in optical flow estimation, even achieving performance comparable to those supervised with ground truth motion on indoor_flying2 and indoor_flying3. This indicates that our design enhances the geometric plausibility of the estimated optical flow while effectively avoiding convergence to local minima, as visualized in Fig. 3.

### v2

@@ Abstract
The estimation of optical flow and 6-DoF ego-motion, two fundamental tasks in 3D vision, has typically been addressed independently. For neuromorphic vision (e.g., event cameras), however, the lack of robust data association makes solving the two problems separately an ill-posed challenge, especially in the absence of supervision via ground truth. Existing works mitigate this ill-posedness by either enforcing the smoothness of the flow field via an explicit variational regularizer or leveraging explicit structure-and-motion priors in the parametrization to improve event alignment. The former notably introduces bias in results and computational overhead, while the latter, which parametrizes the optical flow in terms of the scene depth and the camera motion, often converges to suboptimal local minima. To address these issues, we propose an unsupervised framework that jointly optimizes egomotion and optical flow via implicit spatial-temporal and geometric regularization. First, by modeling camera’s egomotion as a continuous spline and optical flow as an implicit neural representation, our method inherently embeds spatial-temporal coherence through inductive biases. Second, we incorporate structure-and-motion priors through differential geometric constraints, bypassing explicit depth estimation while maintaining rigorous geometric consistency. As a result, our framework (called E-MoFlow) unifies egomotion and optical flow estimation via implicit regularization under a fully unsupervised paradigm. Experiments demonstrate its versatility to general 6-DoF motion scenarios, achieving state-of-the-art performance among unsupervised methods and competitive even with supervised approaches.
@@ 2 Related Work
Event-based Optical Flow Learning
Event cameras asynchronously measure per-pixel intensity changes, not absolute values at fixed intervals [12]. This enables high dynamic range and low-latency sensing. However, estimating optical flow from local event patches inherently suffers from the aperture problem [13, 14]. Because traditional optimization-based methods struggle to recover full motion fields from this ambiguous local data, learning-based approaches have become dominant. These learning-based optical flow methods are broadly categorized as supervised, semi-supervised, and unsupervised.
@@
Event Camera Egomotion Estimation
@@
Flow and Motion Joint Estimation
@@
Neural Flow Fields
Recent studies [21, 49] have demonstrated the significant potential of using deep neural network as continuous and memory-efficient implicit representations for tasks such as view synthesis [50] [51], scene flow estimation [52, 53], and Non-Rigid Structure from Motion [54]. This class of methods leverages coordinate-based networks to store the mapping from spatio-temporal coordinates to target vectors, such as radiance values or motion fields. The inherent smoothness bias of the neural network architecture itself serves as a powerful "neural prior," providing implicit regularization for downstream problems and replacing traditional handcrafted regularizers.
By modeling the differential properties of the scene [52, 53, 54], this paradigm can be extended to represent entire temporal dynamics as Ordinary Differential Equations (ODEs) and to estimate complex, long-term 3D trajectories. This provides a solid foundation for our proposed method.
@@ 4 Experiments
Table 1: Quantitative comparison of optical flow estimation task on MVSEC dataset. Bold is the best among all methods; underlined is second best. Pink represents the best in the ’USL’; Orange represents the second best in the ’USL’.
indoor_flying1 indoor_flying2 indoor_flying3 outdoor_day1 average $dt=1$ EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ %Out $\downarrow$ SL EV-FlowNet-EST [25] 0.97 0.91 1.38 8.20 1.43 6.47 EV-FlowNet+ [29] 0.56 1.00 0.66 1.00 0.59 1.00 0.68 0.99 0.623 0.998 E-RAFT [26] 1.10 5.72 1.94 30.79 1.66 25.20 0.24 1.70 1.235 15.853 DCEIFlow [59] 0.75 1.55 0.90 2.10 0.80 1.77 0.22 0.00 0.668 1.355 TMA [24] 1.06 3.63 1.81 27.29 1.58 23.26 0.25 0.07 1.175 13.563 ADM-Flow [31] 0.52 0.14 0.68 1.18 0.52 0.04 0.41 0.00 0.533 0.340 SSL EV-FlowNet [18] 1.03 2.20 1.72 15.10 1.53 11.90 0.49 0.20 1.193 7.350 STE-FlowNet [33] 0.57 0.10 0.79 1.60 0.72 1.30 0.42 0.00 0.625 0.750 MB Akolkar et al. [13] 1.52 1.59 1.89 2.75 1.938 Nagata et al. [60] 0.62 0.93 0.84 0.77 0.790 Brebion et al. [61] 0.52 0.10 0.98 5.50 0.71 2.10 0.53 0.20 0.685 1.975 Cuadrado et al. [28] 0.58 0.72 0.67 0.85 0.705 Shiba et al. [62] 1.05 2.90 1.68 13.44 1.43 8.97 0.94 3.08 1.275 7.098 MultiCM [17] 0.42 0.09 0.60 0.59 0.50 0.29 0.30 0.11 0.455 0.270 MultiCM-V2 [16] 0.30 0.00 0.47 0.01 0.34 0.00 0.28 0.21 0.348 0.055 USL USL-EV-FlowNet [19] 0.58 0.00 1.02 4.00 0.87 3.00 0.32 0.00 0.698 1.750 FireFlowNet [27] 0.97 2.60 1.67 15.30 1.43 11.00 1.06 6.60 1.283 8.875 ConvGRU-EV-FlowNet [32] 0.60 0.51 1.17 8.06 0.93 5.64 0.47 0.25 0.793 3.615 ET-FlowNeT [35] 0.57 0.53 1.20 8.48 0.95 5.73 0.39 0.12 0.778 3.715 EV-MGRFlowNet [63] 0.41 0.17 0.70 2.35 0.59 1.29 0.28 0.02 0.495 0.958 Paredes et al. [37] 0.44 0.00 0.88 4.51 0.70 2.41 0.27 0.05 0.573 1.743 MotionPriorCMax [36] 0.45 0.09 0.71 2.40 0.60 0.93 Ours 0.40 0.30 0.52 0.18 0.46 0.29 0.42 0.54 0.450 0.328 $dt=4$ SL E-RAFT [26] 2.81 40.25 5.09 64.19 4.46 57.11 0.72 1.12 3.270 40.668 DCEIFlow [59] 2.08 21.47 3.48 42.05 2.51 29.73 0.89 3.19 2.240 24.110 TMA [24] 2.43 29.91 4.32 52.74 3.60 42.02 0.70 1.08 2.762 31.438 ADM-Flow [31] 1.42 7.78 1.88 16.70 1.61 11.40 1.51 10.20 1.605 11.520 SSL EV-FlowNet [18] 2.25 24.70 4.05 45.30 3.45 39.70 1.23 7.30 2.745 29.250 STE-FlowNet [33] 1.77 14.70 2.52 26.10 2.23 22.10 0.99 3.90 1.878 16.700 MB Shiba et al. [62] 4.06 53.88 6.39 71.82 5.36 65.57 3.60 49.04 4.853 60.077 MultiCM [17] 1.68 12.79 2.49 26.31 2.06 18.93 1.25 9.19 1.870 16.805 MultiCM-V2 [16] 1.18 4.77 1.87 15.51 1.38 7.62 1.05 5.68 1.108 8.305 USL USL-EV-FlowNet [19] 2.18 24.20 3.85 46.80 3.18 47.80 1.30 9.70 2.628 32.125 ConvGRU-EV-FlowNet [32] 2.16 21.51 3.90 40.72 3.00 29.60 1.69 12.50 2.688 26.083 ET-FlowNeT [35] 2.08 20.02 3.99 41.33 3.13 31.70 1.47 9.17 2.667 25.555 EV-MGRFlowNet [63] 1.50 8.67 2.39 23.70 2.06 18.00 1.10 6.22 1.763 14.147 Ours 1.58 9.2 2.04 18.54 1.84 13.57 1.63 14.42 1.773 13.933
Table 2: Quantitative comparison of optical flow estimation task on DSEC dataset. Bold is the best among all methods; underlined is second best. Pink represents the best in the ’USL’; Orange represents the second best in the ’USL’.
All interlaken_00_b interlaken_01_a thun_01_a Method EPE $\downarrow$ AE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ AE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ AE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ AE $\downarrow$ %Out $\downarrow$ (SL) E-RAFT [26] 0.79 2.85 2.68 1.39 2.36 6.19 0.90 2.54 3.91 0.65 2.94 1.87 (SL) TMA [24] 0.74 2.68 2.30 1.39 2.16 5.79 0.81 2.23 3.11 0.62 2.88 1.61 (MB) MultiCM-V2 [16] $3.47$ $13.98$ $30.86$ $5.74$ $9.19$ $38.93$ $3.74$ $9.77$ $31.37$ $2.12$ $11.06$ $17.68$ (USL) Paredes et al. [37] 2.33 10.56 17.77 3.34 6.22 25.72 2.49 6.88 19.15 1.73 $9.75$ 10.39 (USL) MotionPriorCMax [36] $3.20$ 8.53 15.21 3.21 4.89 20.45 2.38 5.46 17.40 1.39 6.99 7.36 (USL) EV-FlowNet [19] $3.86$ $31.45$ $6.32$ $47.95$ $4.91$ $36.07$ $2.33$ $20.92$ (USL) Ours 3.14 $10.87$ $19.43$ $7.24$ $14.43$ $35.53$ $3.18$ $7.52$ $19.21$ $1.83$ 6.89 $12.65$ thun_01_b zurich_city_12_a zurich_city_14_c zurich_city_15_a Method EPE $\downarrow$ AE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ AE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ AE $\downarrow$ %Out $\downarrow$ EPE $\downarrow$ AE $\downarrow$ %Out $\downarrow$ (SL) E-RAFT [26] 0.58 2.20 1.52 0.61 4.50 1.06 0.71 3.43 1.91 0.59 2.55 1.30 (SL) TMA [24] 0.55 2.10 1.31 0.57 4.38 0.87 0.66 3.09 1.99 0.55 2.51 1.08 (MB) MultiCM-V2 [17] $2.48$ $12.05$ $23.56$ $3.86$ $28.61$ $43.96$ $2.72$ $12.62$ $30.53$ $2.35$ $11.82$ $20.99$ (USL) Paredes et al. [37] 1.66 $8.41$ 9.34 2.72 23.16 $26.65$ $2.64$ $10.23$ $23.01$ $1.69$ $8.88$ $9.98$ (USL) MotionPriorCMax [36] 1.54 6.55 9.69 $8.33$ 20.16 22.39 1.78 8.79 12.99 1.45 6.27 8.34 (USL) EV-FlowNet [19] $3.04$ $25.41$ 2.62 25.80 $3.36$ $36.34$ $2.97$ $25.53$ (USL) Ours 1.66 5.62 $10.55$ $3.52$ $24.92$ $28.51$ 1.89 6.13 15.44 1.51 5.66 9.08
@@ 4.1 Experimental Setup
We conduct comprehensive evaluations on the MVSEC dataset [64], which is the de facto standard dataset used in prior work to benchmark optical flow and 6-DoF ego-motion estimations. This dataset contains both indoor sequences recorded by drones and outdoor sequences recorded by vehicles. Additionally, we benchmarked optical flow estimation on the more challenging DSEC [65] dataset, which features complex textures, diverse motion patterns, and varying lighting conditions. For optical flow evaluation, we compute three standard metrics: endpoint error (EPE), angular error (AE) and the percentage of pixels with EPE >> 3 pixels (denoted by "% Out"), exclusively over valid ground truth regions with event activity in the evaluation interval. For motion estimation accuracy, we adopt the root mean square error (RMSE) metric proposed in [14] to measure angular velocity ($^{\circ}/s$) and linear velocity ($m/s$) errors.
@@
Optical flow estimation using event camera is a widely explored task with various methodological paradigms. We categorize existing approaches into four primary classes: Supervised Learning(SL), which requires ground truch optical flow for training (e.g., E-RAFT [26], EV-FlowNet-EST [25], EV-FlowNet+ [29], DCEIFlow [59], TMA [24], ADM-Flow [31]); Semi-Supervised Learning (SSL), leveraging grayscale images as supervision through photometric loss construction (e.g., EV-FlowNet [18], STE-FlowNet [33]); Unsupervised Learning (USL), relying solely on event data by warping events to reference time and maximizing accumulated image contrast (e.g., MotionPriorCMax [36], ConvGRU-EV-FlowNet [32], FireFlowNet [27], USL-EV-FlowNet [19], ET-FlowNeT [35], EV-MGRFlowNet [63], Paredes et al. [37]); and Model-Based (MB) methods, which also adopt contrast maximization objectives but employ traditional nonlinear optimization instead of neural networks (e.g., MultiCM-V2[16], MultiCM [17], Akolkar et al. [13], Nagata et al. [60], Brebion et al. [61], Cuadrado et al. [28], Shiba et al. [62]). From these paradigms, we select representative prior works as strong baselines for comprehensive quantitative benchmarking of optical flow estimation performance.
indoor_flying1 indoor_flying2 indoor_flying3 outdoor_day1 average $\text{RMS}_{\omega}\downarrow$ $\text{RMS}_{v}\downarrow$ $\text{RMS}_{\omega}\downarrow$ $\text{RMS}_{v}\downarrow$ $\text{RMS}_{\omega}\downarrow$ $\text{RMS}_{v}\downarrow$ $\text{RMS}_{\omega}\downarrow$ $\text{RMS}_{v}\downarrow$ $\text{RMS}_{\omega}\downarrow$ $\text{RMS}_{v}\downarrow$ w/ D AEmin [42] 1.38 0.069 4.00 0.677 IncEmin [66] 1.65 0.085 4.05 1.035 PEME [43] 1.05 0.039 2.72 0.598 w/o D ECN [48] 0.70 MultiCM-V2 [16] 7.72 0.24 11.50 0.27 9.53 0.31 6.85 5.90 8.900 1.680 Ours 3.44 0.11 5.31 0.12 4.12 0.15 3.38 0.76 4.062 0.285
For 6-DoF motion estimation, existing methods can be categorized into two classes based on whether depth prior knowledge is required. Approaches such as AEmin [42], Incmin [66], and PEME [43] require depth-augmented event data to achieve a 6-DoF estimation, while ECN [48] and MultiCM-V2 [16] can perform a 6-DoF estimation relying solely on raw event streams without depth priors.
@@
The neural implicit flow field adopts the MLP architecture. The detailed network architecture can be found in appendix. Following prior works [17] [43], we partition the entire event sequence into multiple segments during training, with each segment containing 30k events for MVSEC [64] and 300k events for DSEC [65]. Because the time interval of each segment is short, the cubic spline modeling continuous camera motion employs only 4 control knots, whose 6-dimensional vectors are initialized to a constant value of 0.2. To solve the event warping trajectory Eq. (2), we employ the euler solver for its computational efficiency in numerical integration. The weight of the differential geometric loss is set to 0.25 and the differential flow loss to 1. We employ two separate AdamW optimizers [67] for the neural implicit flow field $\text{NN}_{\theta}$ and the camera motion parameters $\beta$. For the MVSEC dataset, the learning rate for the flow field is exponentially decayed from $1\times 10^{-4}$ to $6.3\times 10^{-5}$, whereas for the DSEC dataset, it is cosine annealed from $2\times 10^{-3}$ to $1\times 10^{-7}$. The learning rate for the camera motion is kept constant at $1\times 10^{-3}$ for both datasets. Each short event segment is trained for 1k iterations. All experiments were conducted on a NVIDIA RTX 4090.